### v1

@@ 4.1.3 Evaluation metrics
4.2 Main Results
Table 2: Comparisons of DiffMatch with other semi-supervised methods on LEVIR-CD and WHU-CD testing set with $IoU^{c}$ (%) $\uparrow$ metric. All methods are trained on the classic settings, i.e., the labeled images are selected from the original training sets, which consist of 7,120 samples and 5,947 samples, respectively.
Method Backbone LEVIR-CD(labeled) WHU-CD(labeled)
SemiCDNet (Peng et al., 2020) UNet++ 67.6 71.5 74.3 75.5 51.7 62.0 66.7 75.9
SemiCD (Bandara and Patel, 2022a) ResNet50 72.5 75.5 76.2 77.2 65.8 68.1 74.8 77.2
BAN (Li et al., 2024a) ResNet50 71.3 75.4 80.3 81.6 59.2 64.9 81.1 85.3
SemiVL (Hoyer et al., 2023) ViT-B/16 71.6 74.3 - - - - - -
AdvNet (Vu et al., 2019) ResNet101 66.1 72.3 74.6 75.0 55.1 61.6 73.8 76.6
s4GAN (Mittal et al., 2019) ResNet101 64.0 67.0 73.4 75.4 18.3 62.6 70.8 76.4
Sup. only |||| PSPNet ResNet50 67.5 73.4 75.2 77.7 48.3 60.7 69.7 69.5
UniMatch |||| PSPNet (Yang et al., 2023) ResNet50 75.6 79.0 79.0 78.2 77.5 78.9 82.9 84.4
UniMatch |||| DeepLabV3 (Yang et al., 2023) ResNet50 80.7 82.0 81.7 82.1 80.2 81.7 81.7 85.1
FixMatch (Sohn et al., 2020) ResNet50 79.5 81.3 81.9 81.9 76.5 80.6 81.0 81.8
Table 3: Comparisons of DiffMatch with supervised CD methods.
@@ 4.2 Main Results
FC-EF (Daudt et al., 2018) 83.4 71.5 69.4 53.1
FC-Siam-Diff (Daudt et al., 2018) 86.3 75.9 58.8 41.7
FC-Siam-Conc (Daudt et al., 2018) 83.7 72.0 66.6 50.0
STANet (Chen and Shi, 2020) 87.3 77.4 82.3 70.0
SNUNet (Fang et al., 2021) 88.2 78.8 83.5 71.7
BiT (Chen et al., 2021) 89.3 80.7 84.0 72.4
ChangeFormer (Bandara and Patel, 2022b) 90.4 82.5 89.9 81.6
@@
We select several comparison methods from different perspectives, as listed in Table 2. SemiCDNet (Peng et al., 2020) and SemiCD (Bandara and Patel, 2022a) are specialized SSL models for RSCD. SemiCDNet is built under the adversarial SSL framework, which contains two discriminative networks, one to determine whether the segmentation output is either from unlabeled images or from the ground truth, and the other to encourage similarity between the entropy maps of unlabeled and labeled samples. SemiCD is a consistency regularity-based method that builds learning of unlabeled samples by adding random feature perturbations to the encoder’s difference features and forcing their predictions to be consistent. Another worthy comparison is BAN (Li et al., 2024a), which is a parameter-efficient fine-tuning RSCD framework based on the foundation model, rather than an SSL method, but with excellent transfer capabilities. When the labeled samples are relatively rich (20%, 40%), BAN shows superiority over the previous two RSCD semi-supervised methods, but its performance drops drastically when the samples are fewer (5%, 10%), which illustrates the necessity of learning on unlabeled samples.
SemiVL (Hoyer et al., 2023) is a SOTA semi-supervised segmentation model based on VLM guidance and can be seen as one of the baselines for DiffMatch. We modify SemiVL to fit the CD task, however, it does not show satisfactory performance. We speculate that there are three reasons: 1) the drastic down-sampling in ViT backbones leads to the inability to achieve fine segmentation in RSCD, 2) the lack of a reasonable CEG strategy, and 3) the lack of elaborate design for the CD task.
In addition, some common semi-supervised semantic segmentation models are modified for CD tasks here, in brief, their backbones are reconstructed to siamese structures. AdvNet (Vu et al., 2019) and s4GAN (Mittal et al., 2019) are both adversarial-based models, the former directly or indirectly makes the information entropy of unlabeled data lower, and the latter expects a more similar distribution between labeled and unlabeled samples. FixMatch (Sohn et al., 2020) combines consistency regularization and pseudo labeling while vastly simplifying the overall approach, which we have described in detail in Section 3.1. UniMatch (Yang et al., 2023) adds additional streams of image-level perturbations and feature-level perturbations to FixMatch and supervises them simultaneously using pseudo labels generated by the weak perturbations. On RSCD datasets, FixMatch and UniMatch show decent performance, superior to adversarial-based models. Our method, starting from FixMatch, obtains significant performance gains and achieves optimal $\textit{IoU}^{c}$ score. Specifically, on the LEVIR-CD dataset, 81.9% $\textit{IoU}^{c}$ is achieved when using only 5% labeled data. The improvement on the WHU-CD dataset is more obvious, where the $\textit{IoU}^{c}$ score improves from 76.5% to 81.8%, a span of 5.3%, when using 5% labeled data for training. And the improvement is persistent as the labeled data volume increases. Notably, the components in DiffMatch do not conflict with UniMatch, and their combination can yield further enhancements, but to make DiffMatch clearer, we do not elaborate further coupling with other methods.
In Table 3, we list some recent supervised CD models that are trained on the entire training set. It can be observed that when using only 5% of the labeled data, DiffMatch’s performance is close to that of the models trained with all labels. While using 10% of the labeled data, DiffMatch outperforms the best supervised CD models.
4.3 Cross-dataset Results
Table 4: Comparisons of DiffMatch and other methods under cross-dataset condition. All methods are trained on 5% or 10% labeled data from the LEVIR-CD dataset and unlabeled data from the WHU-CD dataset, and the reported results are evaluated on the LEVIR-CD testing set.
@@ 4.3 Cross-dataset Results
AdvNet (Vu et al., 2019) 65.2 71.6
s4GAN (Mittal et al., 2019) 66.7 68.4
SemiCDNet (Peng et al., 2020) 67.6 71.9
SemiCD (Bandara and Patel, 2022a) 71.4 74.6
FixMatch (Sohn et al., 2020) 75.1 78.3
UniMatch (Yang et al., 2023) 77.0 78.3
@@
To verify whether DiffMatch has the ability to improve CD by utilizing unlabeled data from other datasets, we conduct cross-dataset experiments on LEVIR-CD and WHU-CD datasets. Specifically, we select some samples from the LEVIR-CD dataset as labeled data and images from the WHU-CD dataset as unlabeled data, and evaluate models on the testing set of LEVIR-CD, as listed in Table 4. Compared to our baseline model FixMatch, DiffMatch gets 3.0% and 2.0% improvement to achieve 78.and 80.3% $\textit{IoU}^{c}$ scores when using 5% and 10% labeled data. This experiment demonstrates the potential of DiffMatch to learn from infinite cross-domain data.
4.4 Un-supervised Results
Table 5: Comparisons of our instance-level CEG with other un-supervised CD methods on LEVIR-CD and WHU-CD datasets.
@@ 4.4 Un-supervised Results
PCA-KM (Celik, 2009) 4.8 9.1 5.4 10.2
CNN-CD (El Amin et al., 2016) 7.0 13.1 4.9 9.4
DSFA (Du et al., 2019) 4.3 8.2 4.1 7.8
DCVA (Saha et al., 2019) 7.6 14.1 10.9 19.6
GMCD (Tang et al., 2021) 6.1 11.6 10.9 19.7
DINOv2+CVA (Zheng et al., 2024) - 17.3 - -
AnyChange-H (Zheng et al., 2024) - 23.0 - -
SCM (Tan et al., 2023) 18.8 31.7 18.6 31.3
@@
Table 6: Comparison of pixel-level CEG and instance-level CEG on LEVIR-CD dataset. “Total” denotes the evaluation metrics over all pixels, “Valid” denotes the evaluation metrics over reliable pixels only, i.e., when the probability is greater than the threshold $\gamma$, and “ratio” denotes the ratio of reliable pixels.
Pixel-level Instance-level Total Valid
@@
Our CEG strategy can be directly used as an un-supervised CD method. We compare the instance-level CEG with several traditional and SOTA un-supervised CD methods, as listed in Table 5. Notably, different from earlier methods, DINOv2+CVA (Zheng et al., 2024), AnyChange (Zheng et al., 2024) and SCM (Tan et al., 2023) utilize the power of the foundation model. Specifically, DINOv2+CVA extracts the bi-temporal image embeddings using the pre-trained DINOv2 (Oquab et al., 2023), computes the $l_{2}$ norm of their difference, and generates the binary mask using an adaptive threshold. AnyChange utilizes the capability of the segment anything model (SAM) (Kirillov et al., 2023) to infer instance masks and retrieve their corresponding feature regions, and then computes pixel-level or instance-level feature similarity by cosine distance. However, unfortunately, AnyChange is not semantic-aware since SAM is class-agnostic. Correspondingly, SCM introduces CLIP in the un-supervised CD pipeline to filter out non-specified changes by foreground text guidance. They have achieved a significant improvement over previous methods, benefiting from the accumulated knowledge of the foundation model. Beyond them, our method reaches a new milestone in un-supervised CD, driven by VLM and instance-level CEG. On the testing datasets, our instance-level CEG achieves more than double the improvement over the previous best methods, with $\textit{IoU}^{c}$ improved from 18.8% to 46.3% on the LEVIR-CD dataset and from 18.6% to 45.2% on the WHU-CD dataset. Furthermore, functionally, the instance-level CEG not only generates change masks, but also yields single-temporal semantic masks, which is significant for downstream tasks (e.g., semantic change detection).
In Table 6, we compare pixel-level CEG and instance-level CEG. When using pixel-level CEG, the best total performance is achieved with the threshold $\gamma$ set to 0.5, obtaining 41.4% and 58.5% in $\textit{IoU}^{c}$ and $F_{1}^{c}$ scores. And when using instance-level CEG (threshold $\delta$ defaulted to 0), the $\textit{IoU}^{c}$ and $F_{1}^{c}$ scores are improved to 46.3% and 63.3%, which demonstrates the significant advantage of the instance-level CEG under the purely un-supervised mode.
4.5 Ablation Studies
4.5.1 Effectiveness of Mixed CEG
Table 7: Ablation study of DiffMatch’s components on LEVIR-CD (5% labeled data): vision-language guidance (VL.Guid.), Mixed CEG, dual projection head (DP.Head), decoupled semantic guidance (Dec.Guid.), and contrastive consistency regularization (CCR).
VL.Guid. Mixed CEG DP.Head Dec.Guid. CCR $IoU^{c}$
@@ 4.5.1 Effectiveness of Mixed CEG
Table 8: Ablation study of thresholds (pixel-level) for the generation of pseudo change labels and segmentation labels.
Thres. diff Thres. seg Presicion Recall $IoU^{c}$
@@
Table 9: Ablation study of weights $\lambda_{c}$ and $\lambda_{vl}$ for contrastive loss and VLMguidance loss.
@@
As mentioned in Section 3.2, unlike in the pure un-supervised mode, in the semi-supervised mode we focus more on the precision of the pixels selected to validly provide supervised signals for the unlabeled data, and thus introduce the mixed CEG. In Table 6, we directly evaluate the reliable (valid) pixel portion on the testing set ("Valid" columns). As the pixel-level threshold $\gamma$ is higher, the precision in the reliable pixels improves, which indicates that the pixel threshold $\gamma$ is effective in filtering out the low-precision pixels. However, naturally, the total number of pixels retained decreases as the threshold is increased, implying sparser supervised signals. When both pixel-level CEG and instance-level CEG, i.e., mixed CEG, are used, the highest precision of reliable pixels is achieved, with an improvement in $\textit{IoU}^{c}$ from 65.7% to 69.3%, and the ratio of effective pixels does not decrease, compared to using only pixel-level CEG with a threshold of 0.8.
4.5.2 Components of DiffMatch
@@
Compared to the FixMatch baseline, five components are introduced in DiffMatch, and in Table 7, we analyze them in detail. Additional supervised signals are introduced through the VLM guidance, and the $\textit{IoU}^{c}$ is improved by 1.14% compared to FixMatch. After replacing the pixel-level CEG with the mixed CEG, the performance is improved to 80.97%, which indicates that removing the unaligned pseudo-change regions is effective. The dual projection head avoids the conflict of supervised signals from different sources and further improves the model performance. Notably, the performance decreases after applying CCR, which we attribute to the insufficient guidance to the segmentation decoder. After applying the VLM guidance to the segmentation decoder, the $\textit{IoU}^{c}$ score increases from 81.15% to 81.94%, which again confirms the significance of the VLM guidance. Although five components are added to FixMatch, all of them are activated only during training and do not bring extra reasoning costs.
4.5.3 Hyper-parameter Adjustment
@@
In Table 6, we observe that the performance in un-supervised mode has significant ebb and flow as the pixel-level threshold is changed, so it is essential to explore the effect of these thresholds on DiffMatch in semi-supervised mode. As listed in Table 8, we find that DiffMatch is insensitive to the threshold of the pseudo labels generated by VLM, with only 0.1% undulation in $\textit{IoU}^{c}$ over the threshold change range from 0.5 to 0.9. In addition, an interesting observation is that the model prefers consistent difference threshold and segmentation threshold, which is partially due to CCR, which encourages consistency in segmentation and difference predictions. In fact, this is a side evidence that the CD process is decoupled and that single-temporal scene interpretation and bi-temporal difference detection are endowed with causality.
@@
In Table 9, we conduct ablation studies for the VLM guidance loss $\mathcal{L}_{vl}$ and the contrastive loss $\mathcal{L}_{c}$. It can be observed that both types of losses can bring a certain degree of gain to DiffMatch, however, neither of them can be the major contribution term to the overall loss. Especially for the contrastive loss, when the weight of $\mathcal{L}_{c}$ is set to greater than 0.5, the model performance decreases dramatically. For the VLM guidance loss, this phenomenon exists as well, but is relatively slight. We believe that this is caused by the inevitable presence of errors in the pseudo labels generated by the VLM, which are only capable of performing the auxiliary terms of the CD under the current capabilities of VLM. Further, this explains why we introduce the VLM into the SSL framework instead of direct un-supervised reasoning.
5 Conclusion

### v2

@@ IV-A3 Evaluation metrics
IV-B Main Results
We select several comparison methods from different perspectives, as listed in Table II. SemiCDNet [44] and SemiCD [42] are specialized SSL models for RSCD. SemiCDNet is built under the adversarial SSL framework, which contains two discriminative networks, one to determine whether the segmentation output is either from unlabeled images or from the ground truth, and the other to encourage similarity between the entropy maps of unlabeled and labeled samples. SemiCD is a consistency regularity-based method that builds learning of unlabeled samples by adding random feature perturbations to the encoder’s difference features and forcing their predictions to be consistent. Another worthy comparison is BAN [2], which is a parameter-efficient fine-tuning RSCD framework based on the foundation model, rather than an SSL method, but with excellent transfer capabilities. When the labeled samples are relatively rich (20%, 40%), BAN shows superiority over the previous two RSCD semi-supervised methods, but its performance drops drastically when the samples are fewer (5%, 10%), which illustrates the necessity of learning on unlabeled samples.
SemiVL [28] is a SOTA semi-supervised segmentation model based on VLM guidance and can be seen as one of the baselines for DiffMatch. We modify SemiVL to fit the CD task, however, it does not show satisfactory performance. We speculate that there are three reasons: 1) the drastic down-sampling in ViT backbones leads to the inability to achieve fine segmentation in RSCD, 2) the lack of a reasonable CEG strategy, and 3) the lack of elaborate design for the CD task.
In addition, some common semi-supervised semantic segmentation models are modified for CD tasks here, in brief, their backbones are reconstructed to siamese structures. AdvNet [45] and s4GAN [15] are both adversarial-based models, the former directly or indirectly makes the information entropy of unlabeled data lower, and the latter expects a more similar distribution between labeled and unlabeled samples. FixMatch [18] combines consistency regularization and pseudo labeling while vastly simplifying the overall approach, which we have described in detail in Section III-A. UniMatch [10] adds additional streams of image-level perturbations and feature-level perturbations to FixMatch and supervises them simultaneously using pseudo labels generated by the weak perturbations. On RSCD datasets, FixMatch and UniMatch show decent performance, superior to adversarial-based models. Our method, starting from FixMatch, obtains significant performance gains and achieves optimal $\textit{IoU}^{c}$ score. Specifically, on the LEVIR-CD dataset, 81.9% $\textit{IoU}^{c}$ is achieved when using only 5% labeled data. The improvement on the WHU-CD dataset is more obvious, where the $\textit{IoU}^{c}$ score improves from 76.5% to 81.8%, a span of 5.3%, when using 5% labeled data for training. And the improvement is persistent as the labeled data volume increases. Notably, the components in DiffMatch do not conflict with UniMatch, and their combination can yield further enhancements, but to make DiffMatch clearer, we do not elaborate further coupling with other methods. In Fig. 5, some visual comparisons of the three methods are presented. It is observed that DiffMatch improves in terms of precision and recall, reducing false alarms and missed detections. This implies that the supervision of VLM facilitates the model to distinguish between foreground and background.
TABLE III: Comparisons of DiffMatch with supervised CD methods.
@@ IV-B Main Results
FC-EF [3] 83.4 71.5 69.4 53.1
FC-Siam-Diff [3] 86.3 75.9 58.8 41.7
FC-Siam-Conc [3] 83.7 72.0 66.6 50.0
STANet [20] 87.3 77.4 82.3 70.0
SNUNet [21] 88.2 78.8 83.5 71.7
BiT [46] 89.3 80.7 84.0 72.4
ChangeFormer [47] 90.4 82.5 89.9 81.6
@@
In Table III, we list some supervised CD models that are trained on the entire training set. It can be observed that when using only 5% of the labeled data, DiffMatch’s performance is close to that of the models trained with all labels. While using 10% of the labeled data, DiffMatch outperforms the best supervised CD models. These results demonstrate that our VLM guidance strategy makes the semi-supervised and fully supervised CD methods comparable.
Figure 5: Visualization results of different semi-supervised CD methods on the LEVIR-CD dataset with 5% labels. Red circle indicates the false alarm and yellow circle indicates the missed detection.
IV-C Cross-dataset Results
TABLE IV: Comparisons of DiffMatch and other methods under cross-dataset condition. All methods are trained on 5% or 10% labeled data from the LEVIR-CD dataset and unlabeled data from the WHU-CD dataset, and the reported results are evaluated on the LEVIR-CD testing set.
@@ IV-C Cross-dataset Results
AdvNet [45] 65.2 71.6
s4GAN [15] 66.7 68.4
SemiCDNet [44] 67.6 71.9
SemiCD [42] 71.4 74.6
FixMatch [18] 75.1 78.3
UniMatch [10] 77.0 78.3
@@
To verify whether DiffMatch has the ability to improve CD by utilizing unlabeled data from other datasets, we conduct cross-dataset experiments on LEVIR-CD and WHU-CD datasets. Specifically, we select some samples from the LEVIR-CD dataset as labeled data and images from the WHU-CD dataset as unlabeled data, and evaluate models on the testing set of LEVIR-CD, as listed in Table IV. Compared to our baseline model FixMatch, DiffMatch gets 3.0% and 2.0% improvement to achieve 78.and 80.3% $\textit{IoU}^{c}$ scores when using 5% and 10% labeled data. This experiment demonstrates the potential of DiffMatch to learn from infinite cross-domain data.
IV-D Un-supervised Results
Our CEG strategy can be directly used as an un-supervised CD method. We compare the instance-level CEG with several traditional and SOTA un-supervised CD methods, as listed in Table V. Notably, different from earlier methods, DINOv2+CVA [48], AnyChange [48] and SCM [49] utilize the power of the foundation model. Specifically, DINOv2+CVA extracts the bi-temporal image embeddings using the pre-trained DINOv2 [50], computes the $l_{2}$ norm of their difference, and generates the binary mask using an adaptive threshold. AnyChange utilizes the capability of the segment anything model (SAM) [51] to infer instance masks and retrieve their corresponding feature regions, and then computes pixel-level or instance-level feature similarity by cosine distance. However, unfortunately, AnyChange is not semantic-aware since SAM is class-agnostic. Correspondingly, SCM introduces CLIP in the un-supervised CD pipeline to filter out non-specified changes by foreground text guidance. They have achieved a significant improvement over previous methods, benefiting from the accumulated knowledge of the foundation model. Beyond them, our method reaches a new milestone in un-supervised CD, driven by VLM and instance-level CEG. On the testing datasets, our instance-level CEG achieves more than double the improvement over the previous best methods, with $\textit{IoU}^{c}$ improved from 18.8% to 46.3% on the LEVIR-CD dataset and from 18.6% to 45.2% on the WHU-CD dataset. Furthermore, functionally, the instance-level CEG not only generates change masks, but also yields single-temporal semantic masks, which is significant for downstream tasks (e.g., semantic change detection).
TABLE V: Comparisons of our instance-level CEG with other un-supervised CD methods on LEVIR-CD and WHU-CD datasets.
@@ IV-D Un-supervised Results
PCA-KM [52] 4.8 9.1 5.4 10.2
CNN-CD [53] 7.0 13.1 4.9 9.4
DSFA [54] 4.3 8.2 4.1 7.8
DCVA [55] 7.6 14.1 10.9 19.6
GMCD [56] 6.1 11.6 10.9 19.7
DINOv2+CVA [48] - 17.3 - -
AnyChange-H [48] - 23.0 - -
SCM [49] 18.8 31.7 18.6 31.3
@@
TABLE VI: Comparison of pixel-level CEG and instance-level CEG on LEVIR-CD dataset. “Total” denotes the evaluation metrics over all pixels, “Valid” denotes the evaluation metrics over reliable pixels only, i.e., when the probability is greater than the threshold $\gamma$, and “ratio” denotes the ratio of reliable pixels.
Pixel-level ($\gamma$) Instance-level Total Valid
@@
In Table VI, we compare pixel-level CEG and instance-level CEG. When using pixel-level CEG, the best total performance is achieved with the threshold $\gamma$ set to 0.5, obtaining 41.4% and 58.5% in $\textit{IoU}^{c}$ and $F_{1}^{c}$ scores. And when using instance-level CEG (threshold $\delta$ defaulted to 0), the $\textit{IoU}^{c}$ and $F_{1}^{c}$ scores are improved to 46.3% and 63.3%, which demonstrates the significant advantage of the instance-level CEG under the purely un-supervised mode.
IV-E Ablation Studies
IV-E1 Effectiveness of Mixed CEG
TABLE VII: Ablation study of DiffMatch’s components on LEVIR-CD (5% labeled data): vision-language model guidance (VLM.Guid.), Mixed CEG, dual projection head (DP.Head), decoupled semantic guidance (Dec.Guid.), and contrastive consistency regularization (CCR).
VLM.Guid. Mixed CEG DP.Head Dec.Guid. CCR $IoU^{c}$
@@ IV-E1 Effectiveness of Mixed CEG
As mentioned in Section III-B, unlike in the pure un-supervised mode, in the semi-supervised mode we focus more on the precision of the pixels selected to validly provide supervised signals for the unlabeled data, and thus introduce the mixed CEG. In Table VI, we directly evaluate the reliable (valid) pixel portion on the testing set (”Valid” columns). As the pixel-level threshold $\gamma$ is higher, the precision in the reliable pixels improves, which indicates that the pixel threshold $\gamma$ is effective in filtering out the low-precision pixels. However, naturally, the total number of pixels retained decreases as the threshold is increased, implying sparser supervised signals. When both pixel-level CEG and instance-level CEG, i.e., mixed CEG, are used, the highest precision of reliable pixels is achieved, with an improvement in $\textit{IoU}^{c}$ from 65.7% to 69.3%, and the ratio of effective pixels does not decrease, compared to using only pixel-level CEG with a threshold of 0.8.
IV-E2 Components of DiffMatch
TABLE VIII: Ablation study of thresholds (pixel-level) for the generation of pseudo change labels and segmentation labels.
$\gamma$ $\beta$ Presicion Recall $IoU^{c}$
@@ IV-E2 Components of DiffMatch
TABLE IX: Ablation study of weights $\lambda_{c}$ and $\lambda_{vl}$ for contrastive loss and VLMguidance loss.
@@
Compared to the FixMatch baseline, five components are introduced in DiffMatch, and in Table VII, we analyze them in detail. Additional supervised signals are introduced through the VLM guidance, and the $\textit{IoU}^{c}$ is improved by 1.14% compared to FixMatch. After replacing the pixel-level CEG with the mixed CEG, the performance is improved to 80.97%, which indicates that removing the unaligned pseudo-change regions is effective. The dual projection head avoids the conflict of supervised signals from different sources and further improves the model performance. Notably, the performance decreases after applying CCR, which we attribute to the insufficient guidance to the segmentation decoder. After applying the VLM guidance to the segmentation decoder, the $\textit{IoU}^{c}$ score increases from 81.15% to 81.94%, which again confirms the significance of the VLM guidance. Although five components are added to FixMatch, all of them are activated only during training and do not bring extra reasoning costs.
IV-E3 Hyper-parameter Adjustment
@@
In Table VI, we observe that the performance in un-supervised mode has significant ebb and flow as the pixel-level threshold is changed, so it is essential to explore the effect of these thresholds on DiffMatch in semi-supervised mode. As listed in Table VIII, we find that DiffMatch is insensitive to the threshold of the pseudo labels generated by VLM, with only 0.1% undulation in $\textit{IoU}^{c}$ over the threshold change range from 0.5 to 0.9. In addition, an interesting observation is that the model prefers consistent CD threshold $\gamma$ and segmentation threshold $\beta$, partially due to CCR, which encourages consistency in segmentation and CD predictions. In fact, this is a side evidence that the CD process is decoupled and that single-temporal scene interpretation and bi-temporal change detection are endowed with causality.
@@
In Table IX, we conduct ablation studies for the VLM guidance loss $\mathcal{L}_{vl}$ and the contrastive loss $\mathcal{L}_{c}$. It can be observed that both types of losses can bring a certain degree of gain to DiffMatch, however, neither of them can be the major contribution term to the overall loss. Especially for the contrastive loss, when the weight of $\mathcal{L}_{c}$ is set to greater than 0.5, the model performance decreases dramatically. For the VLM guidance loss, this phenomenon exists as well, but is relatively slight. We believe that this is caused by the inevitable presence of errors in the pseudo labels generated by the VLM, which are only capable of performing the auxiliary terms of the CD under the current capabilities of VLM. Further, this explains why we introduce the VLM into the SSL framework instead of direct un-supervised reasoning.
V Conclusion